A die is thrown 10 times. The probability of getting a three on any throw is 1/6. What is the probability that in the 10 throws, exactly 2 of them are the number three?
Answer #1

Binomial Distribution

n = 10

p = 1/6

q = 1- p = 5/6

So,

Answer is:

0.2907